### v2

@@ 4.3 Specific cases
We now verify (34) by combining this with the ($\Gamma_{G}$-strong) subdifferentiability of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$.
@@
We now interpret the result for standard “scalar” forward backward splitting, as was as outer forward backward splitting for bilevel problems with inner and adjoint methods satisfying the tracking theory of Section 2.
Example 4.13 (Standard forward-backward splitting on a Hilbert space).
Let $X$ be a Hilbert space, and suppose $\nabla F$ is $L$-Lipschitz for some $L\geq 0$, Pick a step length parameter $\tau>0$. The exact forward-backward splitting
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\nabla F(x^{k})$
trivially satisfies the assumptions of the tracking theory in Section 2.1 with $\widetilde{F^{\prime}}\equiv F^{\prime}$, $S_{u}\equiv 0$, $S_{w}\equiv 0$, $\kappa_{u}=\kappa_{w}=1$, $\pi_{u}=\pi_{w}=\mu_{u}=0$ as well sa $\alpha_{u}=\alpha_{w}=0$. It follows that $e_{p,k}\equiv 0$, hence $\varepsilon_{\mathrm{desc},k}\equiv 0$ and $\varepsilon_{k}(\bar{x})\equiv 0$.
To model this algorithm in the operator-relative framework of Theorem 4.11, we take $M=\tau^{-1}\operatorname{Id}$ and $\Lambda=L\operatorname{Id}$. Then the condition $\breve{\Lambda}\leq 2(1-\eta)M$ with $\eta>0$ in Theorem 4.11Item (i) reduces to the standard step length condition $\tau L<2$.
With $G$ $\gamma_{G}$-strongly subdifferentiable for some $\gamma_{G}\geq 0$, and $F^{\prime}$ locally $\gamma_{F}$-monotone for some $\gamma_{F}\in\mathbb{R}$, taking $\Gamma_{G}=\gamma_{G}\operatorname{Id}$, $\Gamma_{F}=\gamma_{F}\operatorname{Id}$, and $\breve{\Lambda}=\breve{\lambda}\operatorname{Id}$ for some $\breve{\lambda}\geq 0$, the conditions in Item 4 reduce to finding $\tau,\beta>0$ and $\eta\geq 0$ such that
$0\leq\tau[L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]\leq 1-\eta\quad\text{and}\quad 0\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|].$
The conditions of Item 3 are analogous.
@@
Example 4.14 (Single-loop/inexact outer forward-backward splitting on a Hilbert space).
Let $X$ be a Hilbert space, and suppose $\nabla F$ is $L$-Lipschitz for some $L\geq 0$. Construct $\widetilde{\nabla F}$ following Section 2.1, for example, using the inner methods of Examples 2.2, 2.4 and 2.5 and the adjoint methods and differential transformation of Examples 2.6 and 2.7. For a step length parameter $\tau>0$, consider the inexact forward-backward splitting
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\widetilde{\nabla F}(x^{k}))$
In the operator-relative framework of Theorem 4.11, we again take $M=\tau^{-1}\operatorname{Id}$ and $\Lambda=L\operatorname{Id}$. Then the condition in Theorem 4.11Item (i) reduces to $20\leq\tilde{\gamma}+\tau(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L<$ 2, where we again recall that $\varsigma_{p}$ is defined in (14).
Suppose then that $G$ is $\gamma_{G}$-strongly subdifferentiable for some $\gamma_{G}\geq 0$, and $F^{\prime}$ is locally $\gamma_{F}$-monotone for some $\gamma_{F}\in\mathbb{R}$. Then the conditions in Item 4 reduce to the bounds
$0\leq\tau[(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\gamma_{F}|-\gamma_{F})]\leq 1-\eta\quad\text{and}\quad 0\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|]-\tilde{\gamma}$
@@
holding for some $\beta,\tilde{\gamma}>0$ and $\eta\geq 0$. Similarly we can translate Item 3.
@@
Thus, we can expect the subdifferential convergence Section 4.2 to hold for small enough $\tau>0$, but stronger forms of convergence via Section 4.2 will require the differential approximation to be good enough that $\varsigma_{p}^{2}L$ is small with respect to the available strong monotonicity. Studying (14), the main constants that control this quantity are $\pi_{u}$ and $\pi_{w}$, which can always be made small if, instead a single iteration, we take sufficiently many iterations of the inner and adjoint solvers that satisfy Section 2.1Items (i) and (ii).
@@
We can also translate Theorem 4.11 to the outer PDPS of Example 4.2 when inner and adjoint methods for estimating the smooth component $f$ of the problem (29) satisfy the tracking theory of Section 2. This is the content of the next theorem. If $f^{\prime}$ is locally $\gamma_{f}$-strongly monotone, its conditions Eqs. 40 and 41 for the satisfaction of Section 4.2 can be ensured by taking the free parameters $\tilde{\gamma},\zeta,\beta>0$ sufficiently small (to ensure Eqs. 40a and 41a), and then $\lambda>0$ sufficiently large (to ensure Eqs. 40b and 41b). Through the basic step length conditions Section 4.1, a large $\lambda$ will force the primal step length $\tau>0$ to small, and allow the dual step length $\sigma>0$ to be correspondingly large. The initial iterate should also satisfy the locality bound (39), where both $\lambda$ and $\tau$ appear, the latter through $M$ defined in (31). For Section 4.2 only the basic step length conditions are required besides the tracking estimates on $f$. Because $\Xi$ is not cyclically monotone (see [26, Chapter 24]), we see no way in general for the PDPS to satisfy Section 4.2Items (iii) and (iv). [^1] However, we could try to enforce the conditions, monitoring for convergence failure by setting expected bounds on $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})=[F+G](x^{N})-[F+G](x^{0})-\sum_{k=0}^{N-1}\langle\Xi x^{k+1}|x^{k}\rangle.$ In fact, if $\inf F+G>-\infty$, we only need to ensure that the latter sum term sum stays within chosen bounds, without having to calculate potentially costly function values.
@@
Assume the setup of Example 4.2 with $g$ and $h_{*}$ convex, and $f$ with an $LM_{z}$-firmly Lipschitz Fréchet derivative for a $L\geq 0$. Suppose that Section 2.1 holds for $f$ in $\Omega\subset Z$ with [^2] We again recall that $d_{Z^{*}}(z^{*},\tilde{z}^{*})=\|z^{*}-\tilde{z}^{*}\|_{M^{-1}_{z}}$ if $M_{z}$ is invertible, in particular with $M_{z}=\operatorname{Id}$ for $Z$ a Hilbert space.
$d_{Z}(z,\tilde{z})=\|z-\tilde{z}\|_{M_{z}},\ d_{Z^{*}}(z^{*},\tilde{z}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M_{z}}^{2})^{*}(z^{*}-\tilde{z}^{*})]^{1/2},\ \text{and}\ b_{Z}(z,\tilde{z})=\|x-\tilde{x}\|_{LM_{z}}.$
Ensure the step length Section 4.1 for some $\tau,\sigma,\lambda>0$. Then
@@
    Section 4.2 holds.
Suppose further that $g$ and $h_{*}$ are, respectively, $\gamma_{g}M_{z}$ and $\gamma_{h_{*}}M_{y}$(-strongly) subdifferentiable for some $\gamma_{g},\gamma_{h_{*}}\geq 0$, and that $f^{\prime}$ is $\gamma_{f}M_{z}$-monotone in $\Omega_{\bar{z}}:=\Omega\ni\bar{z}$ for some $\gamma_{f}\geq 0$. Let $\bar{x}\in\{\bar{z}\}\times\operatorname{dom}h_{*}$ and $\Omega_{\bar{x}}:=\Omega_{\bar{z}}\times\operatorname{dom}h_{*}$. Suppose $\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\subset\Omega_{\bar{z}}$ for some $\delta_{z}>0$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. Set $\varepsilon_{k}(\bar{x})=e_{p,k}(z^{k+1})/(2\lambda\tilde{\gamma})$, and define $r_{p}$ by (35). If
@@
    Section 4.2 option Item (a) holds if, for $\tilde{\gamma}_{f}:=\gamma_{f}-(\zeta/2)L$ and some $\beta,\zeta>0$ and $\eta\geq 0$,
@@
    Section 4.2 option Item (b) holds if $\Omega_{\bar{z}}$ is convex, and, for some $\beta>0$ and $\eta\geq 0$,
@@
Proof 4.16.
$F^{\prime}$ is $\Lambda$-firmly Lipschitz and $\Gamma_{F}$-monotone, and $G$ is $\Gamma_{G}$-strongly convex for
@@
Adopting the distances (36) for $X$ and $X^{*}$, Lemma 4.3 shows that
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}((z^{*},0))=\sup_{(z,y)}\langle z^{*}|z\rangle-\frac{1}{2}\|(z,y)\|_{M}^{2}\leq\sup_{z}\langle z^{*}|z\rangle-\frac{\lambda}{2}\|z\|_{M_{z}}^{2}=\frac{1}{\lambda}(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M_{z}}^{2})^{*}(z^{*}).$
Combining this with Section 2.1 for $f$ and $\widetilde{f^{\prime}}$ in $\Omega\subset Z$, we see that Section 2.1 holds in $\Omega\times\operatorname{dom}h_{*}\subset X$ for $F$ and $\widetilde{F^{\prime}}$ defined in Example 4.2 with $\alpha_{u}$ and $\alpha_{w}$ divided by $\lambda$ compared to the case of $f$ and $\widetilde{f^{\prime}}$. This has the effect of dividing $e_{p,k}$ by $\lambda$. Our claims thus follow if we prove the remaining assumptions of Theorem 4.11.
Item (i): Lemma 4.3 proves $\Lambda\leq(L/\lambda)M$. Clearly $\Lambda\geq 0$. Now we use Theorem 4.11Item (ii).
[…]
@@
Taking $\delta:=\lambda\delta_{z}$, (39) implies, as required, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $2r_{p}<\delta^{2}$. By (42), we have $\mathbb{O}_{M}(\bar{x},\delta)\subset\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\times\operatorname{dom}h_{*}\subset\Omega_{\bar{z}}\times\operatorname{dom}h_{*}=\Omega_{\bar{x}}$. By construction and assumption, we have $\breve{\Lambda}\geq 0$. The claim now follows from Theorem 4.11Item 4.

### v3

@@ 6.1 A general result
Similarly to the claim Item˜3, we now verify (28) by combining this with the ($\Gamma_{G}$-strong) subdifferentiability of $G$, the skew-symmetricity of $\Xi$, and the assumption $(\gamma+\tilde{\gamma})M\leq\Gamma_{G}+\Gamma_{F}-\beta|\Gamma_{F}|$.
Remark 6.5 (error term).
Recalling Remark˜2.7, we can make $r_{\mathrm{desc}}\geq 0$ and $r_{p}\geq 0$ arbitrarily small by taking high-quality first inner and adjoint steps.
Remark 6.6 (linear convergence).
From (52) and Theorem˜2.5, we see that through good-quality first inner and adjoint steps, (41) can be made to hold. Then Corollary˜4.12 is applicable to results satisfying Section˜4.2 for $p>1$ through Theorem˜6.3.
6.2 Forward-backward splitting
@@
We now interpret Theorem˜6.3 for both standard exact forward-backward splitting in a Hilbert space, as well as outer forward-backward splitting when we construct $\widetilde{\nabla F}$ with inner and adjoint methods that satisfy the tracking theory of Section˜2; in particular, the methods of Section˜3.
For clarity of the statement of the next corollary, we omit any mention of parameters such as $\eta$ and $\tilde{\eta}$ that are not important for the algorithm itself, and that can be deduced from the other choices.
@@
Corollary 6.7 (Inexact outer forward-backward splitting on a Hilbert space).
On a Hilbert space $X$, suppose $F:X\to\mathbb{R}$ has an $L$-Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Pick a step length parameter $\tau>0$, and for all $k\in\mathbb{N}$, construct $\widetilde{\nabla F}(x^{k})$ obeying Section˜2.1 for an $\Omega\subset X$ and the distances [^3] The factors in these distances, which arise from the convenience of the operator-relative approach of Theorem 6.3, can always be compensated for by the factors $\pi_{u},\pi_{w},\alpha_{u}$, and $\alpha_{w}$ of Section 2.1.
${d_{X}(x,\tilde{x})}=\tau^{-1}\|x-\tilde{x}\|_{X},\enskip{d_{X^{*}}(x^{*},\tilde{x}^{*})}=\tau\|x^{*}-\tilde{x}^{*}\|_{X},\enskip{b_{X}(x^{k+1},x^{k})}=L\|x^{k+1}-x^{k}\|_{X}.$
Update
Let $\varsigma_{p}$, $\kappa$ and $e_{p,k}$ be as in Theorem˜2.5. Then:
  1. (i)
    Section˜4.2 holds provided $\inf[F+G]>$-$\infty$, $0\leq L\tau+2\varsigma_{p}\sqrt{L\tau}<2$, and
[…]
  2. (ii)
    Section˜4.2 holds.
Suppose further that $G$ is $\gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}(\bar{x},\tau\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\gamma_{F},\gamma_{G}\in\mathbb{R}$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If
$x^{0}\in\mathbb{O}\left(\bar{x},\tau\sqrt{\delta^{2}-2r_{p}}\right)\quad\text{with}\quad r_{p}:=\sup_{N\in\mathbb{N}}\sum_{k=0}^{N-1}p^{k-N}e_{p,k}/(2\tilde{\gamma})<\delta^{2}/2,$
then:
  1. 3.
@@
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{F}:=\gamma_{F}-(\zeta/2)L$ and some $\zeta,\beta>0$,
    $(p-1)/2$ $\leq\gamma:=\tau[\gamma_{G}+\tilde{\gamma}_{F}-\beta|\tilde{\gamma}_{F}|]-\tilde{\gamma}\quad\text{and}$
    $0$ $\leq\tau[(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})]<1$
  2. 4.
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$,
    $p-1$ $\leq\gamma:=\tau[\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|]-\tilde{\gamma}\quad\text{and}$
    $0$ $\leq\tau[(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]<1.$
Proof 6.8.
We apply Theorem˜6.3, whose conditions we need to verify. In its operator-relative framework, we take $\Xi=0$, $M=\tau^{-1}\operatorname{Id}$, and $\Lambda=L\operatorname{Id}$. Then the implicit step (49) holds by (53), and the condition $\Lambda\leq cM$ for a $c>0$ of Theorem˜6.3Item˜(ii) reduces to $\tau L\leq c$, which automatically holds. Further, when
  1. 1.
    in Theorem˜6.3Item˜(i), for (the minimising choice) $\tilde{\gamma}:=(\varsigma_{p}^{2}L\tau)^{-1/2}$ and $0\leq\breve{\lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\tilde{\gamma}\tau^{-1}<\tau^{-1}$, we take $\eta=1-\tau\breve{\lambda}/2>0$;
  2. 2.
    in Theorem˜6.3Item˜3 for $0\leq\breve{\lambda}:=(1+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+\beta^{-1}|\gamma_{F}|-\gamma_{F}<\tau^{-1}$ we take $\eta=1-\tau\breve{\lambda}>0$;
  3. 3.
    in Theorem˜6.3Item˜4 for $0\leq\breve{\lambda}:=(\zeta^{-1}+\varsigma_{p}^{2}\tilde{\gamma}^{-1})L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})<\tau^{-1}$ we take $\eta=1-\tau\breve{\lambda}>0$;
and in each case we take $\breve{\Lambda}=\breve{\lambda}\operatorname{Id}$ and $\tilde{\eta}=0$, then the conditions of Theorem˜6.3Items˜(i), 4 and 3 readily translate to the present respective conditions.
Remark 6.9.
@@
According to the corollary, we can expect the subdifferential convergence Section˜4.2 to hold for small enough $\tau>0$, but stronger forms of convergence via Section˜4.2 will require the differential estimate to be good enough that $\varsigma_{p}^{2}L$ is small with respect to the available strong monotonicity. Studying (13), the main constants that control $\varsigma_{p}$ are $\pi_{u}$ and $\pi_{w}$. They can always be reduced if we repeatedly apply the inner and adjoint solvers that satisfy Section˜2.1Items˜(i) and (ii).
For exact forward-backward splitting with $\widetilde{F^{\prime}}(x^{k})=F^{\prime}(x^{k})$, by taking $p=1$, $\varsigma_{p}=0$, and $e_{p,k}=0$, and then letting γ~
0
[…]
in the previous result, we immediately obtain the following corollary. In Item˜(i), we even take $\Omega=X$, since the tracking inequalities now trivially work with that choice, and we do not assume any local properties form $F$ and $G$; in Items˜3 and 4 we do.
Corollary 6.10 (Exact outer forward-backward splitting on a Hilbert space).
On a Hilbert space $X$, suppose $F:X\to\mathbb{R}$ has an $L$-Lipschitz Fréchet derivative, and $G:X\to\overline{\mathbb{R}}$ is convex, proper, and lower semicontinuous. Pick a step length parameter $\tau>0$. Update
$x^{k+1}:=\operatorname{prox}_{\tau G}(x^{k}-\tau\nabla F(x^{k})).$
Then:
  1. (i)
    Section˜4.2 holds provided $\inf[F+G]>$-$\infty$, and $0\leq L\tau<2.$
  2. (ii)
    Section˜4.2 holds.
Suppose further that $G$ is $\gamma_{G}$-strongly subdifferentiable, and $F^{\prime}$ is $\gamma_{F}$-monotone in $\Omega_{\bar{x}}=\Omega\supset\mathbb{O}(\bar{x},\tau\delta)$ for an $\bar{x}\in X$, $\delta>0$, and $\gamma_{F},\gamma_{G}\in\mathbb{R}$. Pick $\tilde{\gamma}>0$ and $p\in[1,\kappa)$. If $x^{0}\in\mathbb{O}(\bar{x},\tau\delta)$, then:
  1. 3.
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{F}:=\gamma_{F}-(\zeta/2)L$ and some $\zeta,\beta>0$, we have $0\leq\gamma_{G}+\tilde{\gamma}_{F}-\beta|\tilde{\gamma}_{F}|$ and $0\leq\tau[\zeta^{-1}L+2(\beta^{-1}|\tilde{\gamma}_{F}|-\tilde{\gamma}_{F})]<1.$
  2. 4.
    Section˜4.2 option Item˜(b) holds if $\Omega$ is convex, and, for some $\beta>0$, we have $0\leq\gamma_{G}+\gamma_{F}-\beta|\gamma_{F}|$ and $0\leq\tau[L+\beta^{-1}|\gamma_{F}|-\gamma_{F}]<1.$
Now that we have provided step length and growth conditions that prove LABEL:, 4.2, 4.2 and 4.2 for both exact and single-loop forward-backward splitting for bilevel problems, we can use Theorems˜4.6, 4.13 and 4.16 to prove convergence.
6.3 Primal-dual proximal splitting
We consider now the problem (23), i.e.,
(54) $\min_{z\in Z}\penalty 10000\ f(z)+g(z)+h(Kz)=\min_{z\in Z}\max_{y\in Y}\penalty 10000\ f(z)+g(z)+\langle y|Kz\rangle_{Y,Y^{*}}-h_{*}(y),$
where $Z$ and $Y$ are normed spaces equipped with self-adjoint and positive semi-definite $M_{z}\in\mathbb{L}(Z;Z^{*})$ and $M_{y}\in\mathbb{L}(Z;Z^{*})$ that work to generate the fundamental (semi-)norms in these spaces, respecting the Pythagoras’ identity (36). In Hilbert spaces, we would simply take $M_{z}=\operatorname{Id}$ and $M_{y}=\operatorname{Id}$. The functions $g:Z\to\overline{\mathbb{R}}$ and $h:Y^{*}\to\overline{\mathbb{R}}$ are convex, proper, and lower semicontinuous, and $f:Z\to\mathbb{R}$ possibly non-convex but Fréchet differentiable with $LM_{z}$-firmly Lipschitz Fréchet derivative for an $L\geq 0$. Moreover, $K\in\mathbb{L}(Z;Y^{*})$ and $h=(h_{*})^{*}$ for some $h_{*}:Y\to\overline{\mathbb{R}}$.
We will translate Theorem˜6.3 to the outer PDPS of Example˜4.1 when we estimate $f^{\prime}(z^{k})$ through the tracking theory of Section˜2. In Hilbert spaces, we would thus be interested in the method (24), i.e.,
$\left\{\begin{array}[]{l}z^{k+1}:=\operatorname{prox}_{\tau g}(z^{k}-\tau\widetilde{\nabla f}(z^{k})-\tau K^{*}y^{k}),\\ y^{k+1}:=\operatorname{prox}_{\sigma h_{*}}(y^{k}+\sigma K(2z^{k+1}-z^{k})).\end{array}\right$.
In normed spaces, we only represent the problem and method in the implicit forms (23) and (25) with
(55) $F(z,y):=f(z),\quad G(z,y):=g(z)+h_{*}(y),\quad\text{and}\quad\Xi:=\begin{pmatrix}0&K^{*}\\ -K&0\end{pmatrix}$
while the inexact PDPS becomes an instance of (49) with
We start by extending the standard step length assumptions of the PDPS into normed spaces. In the next lemma, in the Hilbert space setting with $M_{y}=\operatorname{Id}$ and $M_{z}=\operatorname{Id}$, we can take $K_{z}=K$ and $K_{y}=\operatorname{Id}$.
Assumption 6.10 (PDPS step length condition).
Suppose $K=K_{y}K_{z}$ for some $K_{z}\in\mathbb{L}(Z;V)$, $K_{y}\in\mathbb{L}(V;Y^{*})$, and a normed space $V$. Given $\lambda\geq 0$, the step length parameters $\tau,\sigma>0$ satisfy
Lemma 6.11 (PDPS preconditioning operator).
If Section˜6.3 holds, then $M$ is positive semi-definite and for any $\gamma_{z},\gamma_{y}\geq 0$ and $\gamma:=\min\{\gamma_{z}\tau,\gamma_{y}\sigma\}/2$, we have
Proof 6.12.
By a simple application of Young’s inequality and Section˜6.3, we have
for any $x=(z,y)\in Z\times Y$. This establishes the first claimed inequality. The second follows by using Young’s inequality and Section˜6.3 to establish
@@ 6.3 Primal-dual proximal splitting
We can now translate Theorem˜6.3 to the outer PDPS of Example˜4.1. It is missing the verification of Section˜4.2Items˜(iii) and (iv) for subdifferential convergence. Because $\Xi$ is not cyclically monotone (see [27, Chapter 24]), we see no way in general for the PDPS to satisfy that property. [^4] However, we could try to enforce the conditions, monitoring for convergence failure by setting expected bounds on $\sum_{k=0}^{N-1}\mathcal{G}(x^{k+1};x^{k})=[F+G](x^{N})-[F+G](x^{0})-\sum_{k=0}^{N-1}\langle\Xi x^{k+1}|x^{k}\rangle.$ In fact, if $\inf F+G>$-$\infty$, we only need to ensure that the latter sum term sum stays within chosen bounds, without having to calculate potentially costly function values.
@@
Assume the setup of Eqs.˜54 and 6.3 for some $\tau,\sigma,\lambda>0$. Suppose that Section˜2.1 holds for $f$ in $\Omega\subset Z$ with [^5] We again recall that $d_{Z^{*}}(z^{*},\tilde{z}^{*})=\|z^{*}-\tilde{z}^{*}\|_{M^{-1}_{z}}$ if $M_{z}$ is invertible, in particular with $M_{z}=\operatorname{Id}$ for $Z$ a Hilbert space.
$d_{Z}(z,\tilde{z})=\|z-\tilde{z}\|_{\lambda M_{z}},\ d_{Z^{*}}(z^{*},\tilde{z}^{*})=[2(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{\lambda M_{z}}^{2})^{*}(z^{*}-\tilde{z}^{*})]^{1/2},\ \text{and}\ {b_{Z}(z^{k+1},z^{k})}=\|z^{k+1}-z^{k}\|_{LM_{z}}.$
Then
@@
    Section˜4.2 holds.
[…]
@@
    Section˜4.2 option Item˜(a) holds if, for $\tilde{\gamma}_{f}:=\gamma_{f}-(\zeta/2)L$ and some $\beta,\zeta>0$ and $\eta\geq 0$,
@@
    Section˜4.2 option Item˜(b) holds if $\Omega_{\bar{z}}$ is convex, and, for some $\beta>0$ and $\eta\geq 0$,
@@
Proof 6.14.
Recall the definitions Eqs.˜55 and 56. Observe that $F^{\prime}$ is $\Lambda$-firmly Lipschitz and $\Gamma_{F}$-monotone, and $G$ is $\Gamma_{G}$-strongly convex for
$\Lambda:=\operatorname{diag}(LM_{z},\;0),\quad\Gamma_{F}:=\operatorname{diag}(\gamma_{F}M_{z},\;0),\quad\text{and}\quad\Gamma_{G}:=\operatorname{diag}(\gamma_{G}M_{z},\;\gamma_{h_{*}}M_{y}).$
To use Theorem˜6.3, we directly verify Theorems˜2.5 and 2.8 for $x^{*}=F^{\prime}(x^{k+1})$, instead of proving Section˜2.1 for the extended functions. By Theorem˜2.5 applied to $f$ and $\widetilde{f^{\prime}}$ and Lemma˜6.11, if $\{x^{n}=(z^{n},y^{n})\}_{n=0}^{k}\subset\Omega_{\bar{x}}$, we have for $x^{k+1}=(z^{k+1},y^{k+1})$ and any $\bar{x}=(\bar{z},\bar{y})$ that
$\begin{split}\langle\widetilde{F^{\prime}}(x^{k})-F^{\prime}(x^{k})|x^{k+1}-\bar{x}\rangle_{X^{*},X}&=\langle\widetilde{f^{\prime}}(z^{k})-f^{\prime}(z^{k})|z^{k+1}-\bar{z}\rangle_{Z^{*},Z}\\ &\geq-\frac{\tilde{\gamma}}{2}\|z^{k+1}-\bar{z}\|_{\lambda M_{z}}^{2}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}\|z^{k+1}-z^{k}\|_{LM_{z}}-\frac{1}{2\tilde{\gamma}}e_{p,k}\\ &\geq-\frac{\tilde{\gamma}}{2}\|x^{k+1}-\bar{x}\|_{M}^{2}-\frac{\varsigma_{p}^{2}}{2\tilde{\gamma}}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}-\frac{1}{2\tilde{\gamma}}e_{p,k}.\end{split}$
Recalling the choice of distances (48), this verifies Theorem˜2.5 for $F$ and $\widetilde{F^{\prime}}$ in $\Omega_{\bar{x}}$.
@@
$(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{M}^{2})^{*}((z^{*},0))=\sup_{(z,y)}\langle z^{*}|z\rangle-\frac{1}{2}\|(z,y)\|_{M}^{2}\leq\sup_{z}\langle z^{*}|z\rangle-\frac{\lambda}{2}\|z\|_{M_{z}}^{2}=(\tfrac{1}{2}\|\,\boldsymbol{\cdot}\,\|_{\lambda M_{z}}^{2})^{*}(z^{*}).$
If $\{x^{n}=(z^{n},y^{n})\}_{n=0}^{k}\subset\Omega_{\bar{x}}$, Theorem˜2.8 applied to $f$ and $\widetilde{f^{\prime}}$ now establishes
${d_{X^{*}}^{2}(\widetilde{F^{\prime}}(x^{k}),F^{\prime}(x^{k}))}\leq d_{Z^{*}}^{2}(\widetilde{f^{\prime}}(z^{k}),f^{\prime}(z^{k}))\leq e_{\mathrm{lip},k},$
where $e_{\mathrm{lip},k}$ satisfies (15). The “whenever” of (15) is equivalent to $\sup_{N\in\mathbb{N}}\sum_{j=0}^{N-1}\|x^{k+1}-x^{k}\|_{\Lambda}^{2}<\infty$. This verifies Theorem˜2.8 for $F$ and $\widetilde{F^{\prime}}$.
We proceed with proving our specific claims.
Item˜(i): Lemma˜6.11 proves $\Lambda\leq(L/\lambda)M$. Clearly $\Lambda\geq 0$. Now we use Theorem˜6.3Item˜(ii).
Item˜3: Observe that $|\Gamma_{F}|=\operatorname{diag}(|\gamma_{f}|M_{z},\;0)$. Taking $\gamma_{z}:=\gamma_{g}+\gamma_{f}-\beta|\gamma_{f}|$ and $\gamma_{y}:=\gamma_{h_{*}}$ in Lemma˜6.11, and using Eq.˜59a and (59b), we obtain the required bounds
(60) $(1-\eta)M$ $\geq(1-\eta)\lambda\operatorname{diag}(M_{z},\;0)\geq\operatorname{diag}(\breve{\lambda}M_{z},0)$
@@
Taking $\delta:=\lambda\delta_{z}$, (57) implies, as required, $x^{0}\in\mathbb{O}_{M}(\bar{x},\sqrt{\delta^{2}-2r_{p}})$ and $2r_{p}<\delta^{2}$. By (60), we have $\mathbb{O}_{M}(\bar{x},\delta)\subset\mathbb{O}_{M_{z}}(\bar{z},\delta_{z})\times\operatorname{dom}h_{*}\subset\Omega_{\bar{z}}\times\operatorname{dom}h_{*}=\Omega_{\bar{x}}$. By construction and assumption, we have $\breve{\Lambda}\geq 0$. The claim now follows from Theorem˜6.3Item˜4.